### v1

@@ 2.5 Discovering AGN through optical variability
Given the fact that variability is a prominent feature of AGN emission, it is not surprising that it has been widely used to identify active nuclei in galaxies (see Mushotzky 202, for an introductory review on how to find AGN). This approach was adopted not long after the actual discovery of QSO [307] but became increasingly effective as modern automated photometric techniques were developed at the end of the last century. In fact, the high incidence of variability among quasars observed by [24, 127] led several authors to exploit this property to derive homogeneous samples of sources [e.g. 103]. Focusing on selected sky regions with deep observations and many photometric standard stars, such as the so-called ”Selected Area” (SA) n.57 and n.68, and using observations obtained with the Mayall telescope at Kitty Peak National Observatory, [139] were able to retrieve samples of quasars down to faint magnitudes (e.g. mag) based on a combination of colours, proper motion, and variability measured between two epochs one year apart. Further studies significantly extended the temporal baseline to more than a decade, using Mayall plates, probing down to mag [290, 291].
@@
[274] proposed a different approach to select quasars in large surveys, based on the amplitude and slope of the Structure Function. Using epochs of imaging data, taken from the SDSS Stripe 82 survey over 5 years, they showed that quasars can be effectively separated from other non-variable and variable sources by their location in the plane (Fig. 18). They noted that variability performs just as well as colour selection in the identification of quasars with a completeness of 90% and a purity of 95%, also in redshift ranges where color selection is known to be problematic. Even with much sparser time sampling, e.g. just six epochs over 3 years, variability is an efficient quasar classifier reaching purity and completeness. Similarly, [247] combined and SF parameters derived from Stripe 82 lightcurves but used a supervised Neural Network trained on the BOSS spectroscopic sample, to distinguish stars from quasars. In line with previous findings, they were able to retrieve samples with completeness and about purity, somewhat dependent on the adopted colour cuts. They also observed that Broad Absorption Line (BAL) quasars are preferentially selected by variability than by color criteria. In Palanque-Delabrouille et al. [230, 231] they implemented this approach to study the quasar luminosity function. To further improve this methodology, [178] proposed using a damped random walk model parametrised by the damping timescale and the asymptotic amplitude . In particular, the inclusion of the damping timescale information increased the purity from 60% to 75% while maintaining a highly complete sample (98%) even in the absence of color information. This approach may be limited by the ability of DRW models to reproduce actual AGN lightcurves (see Sect. 2.4) and by the fact that is poorly constrained if the data do not probe timescales significantly longer than the damping one (see Sect. 2.2). Nevertheless, this methodology looks promising for current and next-generation surveys spanning several years (see Sect. 2.6).
@@
Despite all these efforts, probing the faint end of the AGN luminosity function by variability selection has proven to be challenging, because of the contamination of the host galaxy, which strongly dilutes the emission of faint, obscured, and/or high-redshift sources when using ground-based observations. Sarajedini et al. [265, also see ] compared two observations of the Hubble Deep Field, separated by a gap of 5 years, with the goal of investigating the population of active galactic nuclei to . They found evidence of variability in of galaxies down to . About of their AGN candidates are X-ray confirmed, while spectroscopic observations reveal that their sample is mainly composed of type 2 AGN. Furthermore, when stacking the X-ray undetected sources, it is usually found that there is a significant excess of hard X-ray emission (N. Brandt, private communication). On the other hand of the X-ray sources in the field are optically variable, supporting their interpretation as low-luminosity AGN (LLAGN). [52] also studied variable AGN in the Hubble Ultra Deep Field, in the -band, down to 28 mag. They concluded that about 1% of all field objects in the range of show variability. They discussed the possibility of tracing AGN activity induced by merger events through variability, although with somewhat inconclusive results.
@@
This decade will see the beginning of operations of the 8.4 m Vera C. Rubin Observatory888https://rubinobservatory.org, which will conduct the Legacy Survey of Space and Time [LSST, 118] through the 3.2 gigapixel focal array camera. Although the details of the observing strategy are still being finalised [28, also see the LSST website999https://www.lsst.org/content/charge-survey-cadence-optimization-committee-scoc], the LSST will conduct a deep-wide-fast survey that will observe a deg2 region about 800 times (i.e. every few nights in the optical/near-infrared bands, Fig. 23) during the anticipated 10 yr of operations. It will also yield co-added images that will probe astronomical sources down to . These data will include observations of 20 billion galaxies and a similar number of stars, as well as tens of millions of AGN [283]. In addition to the main survey, LSST will also target a number of Deep Drilling Fields (DDF), namely ELAIS-S1, XMM-LSS, CDF-S, COSMOS and Euclid-DFS, which will benefit from a faster cadence resulting in significantly deeper final co-adds [35]. LSST will also take advantage of a number of ancillary surveys, both photometric and spectroscopic, intended to increase the cadence and length of the lightcurves, the wavelength coverage and the spectroscopic completeness [e.g. 278, also see the LSST website101010https://www.lsst.org/scientists/in-kind-program], .
@@
Considering the number of studies and advances triggered by the SDSS Stripe 82 light curves, it is reasonable to assume that LSST will change the field. Preliminary studies have already tried to assess the ability of LSST to detect and characterise AGN through a combination of colours, variability, and proper motions, using simulations and current survey data [e.g. 281, 268, 149, 49, 86, 194], as well as machine learning algorithms [e.g. 71, 138, 83]. Among these efforts, LSST AGN data challenge [269] aimed to validate both classical and machine learning approaches for AGN selection and characterisation, using SDSS Stripe 82 lightcurves as a test ground. The performance of supervised (support vector machine, random forest, extreme gradient boosting, artificial neural network, convolutional neural network) and unsupervised models (deep embedding clustering) obtained a classification accuracy of 97.5% for supervised models and a clustering accuracy of 96.0% for unsupervised ones on a blinded data set. The experiment confirmed that variability features significantly improve the accuracy of the trained models, and correlation analysis among different bands enables a fast and inexpensive first-order selection of quasar candidates. Such attempts only provide a rough estimate of the LSST potential, since they rely on pre-selected, confirmed quasar samples and do not properly probe the average AGN population accessible to the survey. For this reason, other studies, some mentioned before, tested the ability to detect normal AGN either by using image-difference analysis or extended lightcurves probing some of the DDFs down to single-epoch depths comparable to LSST [e.g, 44, 68, 58, 66].
@@
Eventually, even LSST will be hampered by seasonal gaps and seeing limitations and by an expected temporal span of 10 years, which has been proven to be insufficient (although not at the same depth and cadence) to provide a full understanding of optical variability in previous studies (see discussion in Sect. 2.2 and 2.5). For this reason it is essential that the astronomical community embarks in an effort to merge all existing and upcoming data to increase the temporal coverage of many thousand to millions of AGN up to several decades, hopefully allowing the complegte characterisation of the AGN variations and disentangle the processes dominating different temporal frequency domains. In addition, new methods are being developed to deal with sparsely sampled lightcurves and more flexible empirical models are being developed which will help to deal with some of the limitations of current and future datasets [e.g. 162].
@@ 3.1 A brief history of AGN X-ray observations
It was observations made by the Medium Energy (ME; 2-10 keV), and the Low Energy (LE; 0.05-2 keV) instruments on board the EXOSAT observatory that helped us better understand the nature of the X-ray variability of AGN. This was because these instruments were more sensitive than previous X-ray detectors and because of the unique capability of EXOSAT to continuously monitor a source for up to 3 days. Figure 24 shows an example of the EXOSAT “long-look” light curves. AGN appeared to be variable on all the timescales sampled, from minutes to days. The long EXOSAT light curves could also be used for power spectrum analysis, for the first time in the X-ray variability studies of AGN. Since then, power spectral analysis has been the main tool for studying and characterising the X-ray variability of these objects. Thus, before continuing, we briefly summarise the basics of power spectrum analysis, as done in Sect. 2.1 for SF analysis.
@@
where is a complex random function, which determines the amplitude of the sinusoids, , into which we can decompose [see, e.g., 242]. In fact, the amplitude of the sinusoids is proportional to . We therefore define the PSD of the stationary random process as follows:
@@
where is the mean operator, and is usually called “time lag” (or simply “lag”). As discussed also in Sect. 2.1, this function is very useful, as it quantifies the “memory” of the system, which is one of the most important characteristics of a random process. For most of the random processes observed in astrophysics, memory is clearly present in the sense that the value of the process at time depends on its past values. Consequently is non-zero over a wide range of lags. The broader is, the longer the memory of the system. It can be shown that, for a stationary random process, the PSD is the Fourier transform of the autocovariance function of the process, i.e.
The equation above implies that the PSD holds exactly the same information as the auto-covariance does. However, it turns out that the PSD is a much more useful tool to study a variable process because of the statistical properties of the estimators of these functions. Estimators of use the observed points in a light curve and turn out to be highly correlated, exactly due to the intrinsic memory of the process, which one wishes to study. This issue was discussed in Sect. 2.1 as well, when describing the properties of the SF. Consequently, they can be heavily biased towards the intrinsic value of , over a wide range of time-lags. Furthermore, their statistical properties, such as their variance (i.e. their error), are unknown, as they depend on the unknown statistical properties of the variable process under study. Therefore, the sampled autocovariance functions cannot be easily fitted by models.
The most frequently used estimator of the PSD is called the “periodogram”, and its statistical properties are much superior. If computed over a certain range of frequencies, the periodogram estimates are asymptotically independent random variables with a known distribution. Their error depends on the intrinsic PSD (which is unknown), but this deficit can be overcome by considering either the logarithm of the periodogram [in which case the log-periodogram error is know, i.e. 227], or by “smoothing” the periodogram estimates. The periodogram is an asymptotically unbiased estimate of the intrinsic PSD, although in the case of “red-noise” power spectra the lowest frequency estimates can be biased (this is the so-called “red-leakage” bias, see 227, and references therein). Given the superior statistical properties of the PSD estimates and the fact that it can help us easily detect characteristic timescales (either signaling periodicity or otherwise), PSD analysis has been the dominant method of studying the variability in the observed X-ray light curves of AGN.
@@ 3.3 Results from the PSD analysis of the X-ray light curves
Another interesting result from the power spectrum analysis of the EXOSAT light curves was that the ME PSDs were significantly flatter than the lower-energy PSDs [228]. This is in contradiction with the expectations from inverse Compton scattering of low-energy photons in a single-temperature X-ray corona. Hard X-ray photons should undergo more scatterings in this case, hence high frequency variations should be washed out, and the PSD at high energies should be separately steeper, contrary to what is observed. Similar results were also reported for galactic X-ray black hole binaries (GBHBs; e.g. Nowak et al. 213).
@@ 3.3.1 The RXTE/XMM-Newton legacy
It was long proposed that the X-ray PSDs of AGN cannot continue with a steep slope to very low frequencies because the total variance would then be unphysically large. There were indications that PSD “breaks” to a flatter slope at frequencies below a characteristic “break” or “bending” frequency () in the PSD analysis of past observations of AGN [e.g. 229]. However, the first clear and unambiguous detection of such a “break timescale” was based on the PSD analysis of the RXTE light curves of NGC 3516 and was reported by [78]. Soon after, [301] and [184] published the results of a detailed study of the power spectrum of a few Seyferts using RXTE light curves and reported further detections of such bending frequencies. Based on the results of these papers, a picture of a power spectrum with a slope of at low frequencies that steepens to above emerged in the 2–10 keV band.
The next significant improvement in the determination of the AGN X-ray PSDs was made possible by the combination of data from RXTE and XMM-Newton in order to calculate the power spectrum over a very wide frequency range. Arguably, the best example of such a well-defined PSD is the power spectrum of NGC 4051 presented by McHardy et al. [199, see the left panel in Fig. 26]. This power spectrum is well fitted by a “bending power law (BPL) model” of the form:
where and are the PSD slopes below and above , respectively, while is equal to PSD(. In the case where , then PSD at all frequencies below and is representative of the PSD amplitude.
@@
There is plenty of observational evidence on the increase in the characteristic frequencies in the GBHB X-ray power spectra with increasing X-ray flux, that is, with an increasing accretion rate in this case (e.g. Axelsson et al. 4 for Cyg X-1, Reig et al. 255 for GRO J1655-40, Alabarta et al. 3 for MAXI J1727–203). Assuming that the X-ray variability properties of AGN and GBHBs are similar, this observational evidence supports the findings of [193]. If true, this result puts interesting constraints on the physical timescale that could be responsible for in AGN.
@@
It is possible that the corona is highly magnetised and is powered by magnetic processes. For example, according to Haardt et al. [111, see also ] the energy that powers the X-ray corona may be stored in magnetic field structures that produce active X-ray regions on top of the disc. The energy is stored in the magnetic field on a charge timescale, , and is released on a shorter discharge timescale, . In this scenario, the main characteristic timescales of the X-ray corona are and . The charge timescale is given by:
The left panel of Fig. 27 shows a plot of log( versus BH mass, for a small sample of AGN. We used data from [102] for NGC 4395 and [196] for Mrk 766, who studied the power spectra using only XMM-Newton light curves. For all the other objects, we used results from power spectral analysis of combined RXTE and XMM-Newton PSDs (McHardy et al. 199 for NGC 4051, McHardy et al. 187 for MCG -6-30-15, Uttley and McHardy 300 for NGC 3227, Markowitz 173 for NGC 7469, Summons et al. 258 for NGC 3783, and Markowitz et al. 184 for NGC 4151, NGC 3516, NGC 5548 and Fairall 9). For the BH mass estimates, we used the measurements listed in the “AGN Black Hole Mass Database”, which have been calculated by spectroscopic reverberation mapping studies [assuming a mean factor of 4.3, 29]. Solid, dashed, and dotted lines show the theoretical MBH relation assuming that , in the case where and 0.01, respectively. Clearly, the hypothesis that can explain most of the scatter in the observed relation between and MBH, as long as the accretion rate decreases from the smaller to the larger BH mass objects, which is what is observed in low redshift samples of AGN, such as the BASS sample [e.g. 151].
@@
The right panel in Fig. 27 shows the product PSD( for Seyeferts where this product can be calculated using the best-fit results from the BPL model fits to the 2-10 keV PSDs. This quantity is representative of the PSD amplitude (as we explained in §3.3.1). We have used the results of [199, 187, 300] for NGC 4051, MCG-6-30-15, and NGC 3227, respectively, and [184] for Fairall 9, NGC 5548 and NGC 3516. These authors used RXTE and XMM-Newton light curves to compute the power spectrum over a wide range of frequencies. For the rest of the sources (i.e. Mrk 335, Mrk 766, NGC 4395 and NGC 6860) we used the [102] results. PSD amplitudes are plotted as a function of log( (we used the Gupta et al. 99 measurements of to calculate ).
@@
The number of PSD amplitude measurements so far has been rather small, and we cannot draw any firm conclusions. The data so far indicate a distribution of PSD amplitudes with the same mean for all AGN, but further PSD analysis of combined RXTE and XMM light curves is necessary to reach conclusive results.
@@
The PSD slope at frequencies below is generally believed to be very close to . This is a rather well established fact, based on the results of the bending PL model fits to PSDs computed using long-term RXTE light curves. The left panel of Fig. 28 shows the best fit for NGC 4051 [199], MCG-6-30-15 [187], Fairall 9, NGC5548, NGC3516 and NGC4151 [184], as well as NGC3227 [300], NGC3783 [258] and NGC7469 [173]. The solid horizontal line indicates the value of . Clearly, all the low-frequency PSD slope estimates are consistent with this value (within the error). The PSDs of NGC4051 and MCG-6-30-15 (199 and 187, respectively) are particularly important in this respect, as they show that the slope of -1 extends over and orders of magnitude below . This is typical of the XRB PSDs when they are in their so-called “soft state”.

### v3

@@ 2.5 Discovering AGN through optical variability
Given the fact that variability is a prominent feature of AGN emission, it is not surprising that it has been widely used to identify active nuclei in galaxies (see Mushotzky 204, for an introductory review on how to find AGN). This approach was adopted not long after the actual discovery of QSO [309] but became increasingly effective as modern automated photometric techniques were developed at the end of the last century. In fact, the high incidence of variability among quasars observed by [25, 129] led several authors to exploit this property to derive homogeneous samples of sources [e.g. 105]. Focusing on selected sky regions with deep observations and many photometric standard stars, such as the so-called ”Selected Area” (SA) n.57 and n.68, and using observations obtained with the Mayall telescope at Kitty Peak National Observatory, [141] were able to retrieve samples of quasars down to faint magnitudes (e.g. mag) based on a combination of colours, proper motion, and variability measured between two epochs one year apart. Further studies significantly extended the temporal baseline to more than a decade, using Mayall plates, probing down to mag [292, 293].
@@
[276] proposed a different approach to select quasars in large surveys, based on the amplitude and slope of the Structure Function. Using epochs of imaging data, taken from the SDSS Stripe 82 survey over 5 years, they showed that quasars can be effectively separated from other non-variable and variable sources by their location in the plane (Fig. 18). They noted that variability performs just as well as colour selection in the identification of quasars with a completeness of 90% and a purity of 95%, also in redshift ranges where color selection is known to be problematic. Even with much sparser time sampling, e.g. just six epochs over 3 years, variability is an efficient quasar classifier reaching purity and completeness. Similarly, [249] combined and SF parameters derived from Stripe 82 lightcurves but used a supervised Neural Network trained on the BOSS spectroscopic sample, to distinguish stars from quasars. In line with previous findings, they were able to retrieve samples with completeness and about purity, somewhat dependent on the adopted colour cuts. They also observed that Broad Absorption Line (BAL) quasars are preferentially selected by variability than by color criteria. In Palanque-Delabrouille et al. [232, 233] they implemented this approach to study the quasar luminosity function. To further improve this methodology, [180] proposed using a damped random walk model parametrised by the damping timescale and the asymptotic amplitude SF∞. In particular, the inclusion of the damping timescale information increased the purity from 60% to 75% while maintaining a highly complete sample (98%) even in the absence of color information. This approach may be limited by the ability of DRW models to reproduce actual AGN lightcurves (see Sect. 2.4) and by the fact that is poorly constrained if the data do not probe timescales significantly longer than the damping one (see Sect. 2.2). Nevertheless, this methodology looks promising for current and next-generation surveys spanning several years (see Sect. 2.6).
@@
Despite all these efforts, probing the faint end of the AGN luminosity function by variability selection has proven to be challenging, because of the contamination of the host galaxy, which strongly dilutes the emission of faint, obscured, and/or high-redshift sources when using ground-based observations. Sarajedini et al. [267, also see ] compared two observations of the Hubble Deep Field, separated by a gap of 5 years, with the goal of investigating the population of active galactic nuclei to . They found evidence of variability in of galaxies down to . About of their AGN candidates are X-ray confirmed, while spectroscopic observations reveal that their sample is mainly composed of type 2 AGN. Furthermore, when stacking the X-ray undetected sources, it is usually found that there is a significant excess of hard X-ray emission (N. Brandt, private communication). On the other hand of the X-ray sources in the field are optically variable, supporting their interpretation as low-luminosity AGN (LLAGN). [53] also studied variable AGN in the Hubble Ultra Deep Field, in the -band, down to 28 mag. They concluded that about 1% of all field objects in the range of show variability. They discussed the possibility of tracing AGN activity induced by merger events through variability, although with somewhat inconclusive results.
@@
This decade will see the beginning of operations of the 8.4 m Vera C. Rubin Observatory888https://rubinobservatory.org, which will conduct the Legacy Survey of Space and Time [LSST, 120] through the 3.2 gigapixel focal array camera. Although the details of the observing strategy are still being finalised [29, also see the LSST website999https://www.lsst.org/content/charge-survey-cadence-optimization-committee-scoc], the LSST will conduct a deep-wide-fast survey that will observe a deg2 region about 800 times (i.e. every few nights in the optical/near-infrared bands, Fig. 23) during the anticipated 10 yr of operations. It will also yield co-added images that will probe astronomical sources down to . These data will include observations of 20 billion galaxies and a similar number of stars, as well as tens of millions of AGN [285]. In addition to the main survey, LSST will also target a number of Deep Drilling Fields (DDF), namely ELAIS-S1, XMM-LSS, CDF-S, COSMOS and Euclid-DFS, which will benefit from a faster cadence resulting in significantly deeper final co-adds [36]. LSST will also take advantage of a number of ancillary surveys, both photometric and spectroscopic, intended to increase the cadence and length of the lightcurves, the wavelength coverage and the spectroscopic completeness [e.g. 280, also see the LSST website101010https://www.lsst.org/scientists/in-kind-program], .
@@
Considering the number of studies and advances triggered by the SDSS Stripe 82 light curves, it is reasonable to assume that LSST will change the field. Preliminary studies have already tried to assess the ability of LSST to detect and characterise AGN through a combination of colours, variability, and proper motions, using simulations and current survey data [e.g. 283, 270, 151, 50, 88, 196], as well as machine learning algorithms [e.g. 73, 140, 85]. Among these efforts, LSST AGN data challenge [271] aimed to validate both classical and machine learning approaches for AGN selection and characterisation, using SDSS Stripe 82 lightcurves as a test ground. The performance of supervised (support vector machine, random forest, extreme gradient boosting, artificial neural network, convolutional neural network) and unsupervised models (deep embedding clustering) obtained a classification accuracy of 97.5% for supervised models and a clustering accuracy of 96.0% for unsupervised ones on a blinded data set. The experiment confirmed that variability features significantly improve the accuracy of the trained models, and correlation analysis among different bands enables a fast and inexpensive first-order selection of quasar candidates. Such attempts only provide a rough estimate of the LSST potential, since they rely on pre-selected, confirmed quasar samples and do not properly probe the average AGN population accessible to the survey. For this reason, other studies, some mentioned before, tested the ability to detect normal AGN either by using image-difference analysis or extended lightcurves probing some of the DDFs down to single-epoch depths comparable to LSST [e.g, 45, 70, 59, 68].
@@
Eventually, even LSST will be hampered by seasonal gaps and seeing limitations and by an expected temporal span of 10 years, which has been proven to be insufficient (although not at the same depth and cadence) to provide a full understanding of optical variability in previous studies (see discussion in Sect. 2.2 and 2.5). For this reason it is essential that the astronomical community embarks in an effort to merge all existing and upcoming data to increase the temporal coverage of many thousand to millions of AGN up to several decades, hopefully allowing the complegte characterisation of the AGN variations and disentangle the processes dominating different temporal frequency domains. In addition, new methods are being developed to deal with sparsely sampled lightcurves and more flexible empirical models are being developed which will help to deal with some of the limitations of current and future datasets [e.g. 164].
@@ 3.1 A brief history of AGN X-ray observations
It was observations made by the Medium Energy (ME; 2-10 keV), and the Low Energy (LE; 0.05-2 keV) instruments on board the EXOSAT observatory that helped us better understand the nature of the X-ray variability of AGN. This was because these instruments were more sensitive than previous X-ray detectors and because of the unique capability of EXOSAT to continuously monitor a source for up to 3 days. Figure 24 shows an example of the EXOSAT “long-look” light curves. AGN appeared to be variable on all the timescales sampled, from minutes to days. The long EXOSAT light curves could also be used for power spectrum analysis, for the first time in the X-ray variability studies of AGN. Since then, power spectral analysis has been the main tool for studying and characterising the X-ray variability of these objects. Thus, before continuing, we briefly summarise the basics of power spectrum analysis, as done in Sect. 2.1 for the structure function analysis.
@@
where is a complex random function, which determines the amplitude of the sinusoids, , into which we can decompose [see, e.g., 244]. In fact, the amplitude of the sinusoids is proportional to . We therefore define the PSD of the stationary random process as follows:
@@
where is the mean operator, is the mean of the process, and is usually called “time lag” (or simply “lag”). As discussed also in Sect. 2.1, this function is very useful, as it quantifies the “memory” of the system, which is one of the most important characteristics of a random process. For most of the random processes observed in astrophysics, memory is clearly present in the sense that the value of the process at time depends on its past values. Consequently is non-zero over a wide range of lags. The broader is, the longer the memory of the system. It can be shown that, for a stationary random process, the PSD is the Fourier transform of the autocovariance function of the process, i.e.
The equation above implies that the PSD holds exactly the same information as the auto-covariance does. We can express as the inverse Fourier transform of . Since , we can then write
This equation highlights another useful property of the power spectral density function. The product is the contribution to the total variance of the random process of sinusoids in with frequencies between and . Furthermore, it turns out that PSD is a more useful tool to study a variable process due to the statistical properties of the estimators of these functions. Estimators of use the observed points in a light curve and turn out to be highly correlated, exactly due to the intrinsic memory of the process, which one wishes to study. This issue was discussed in Sect. 2.1 as well, when describing the properties of the SF. Consequently, they can be heavily biased towards the intrinsic value of , over a wide range of time-lags. Furthermore, their statistical properties, such as their variance (i.e. their error), are unknown, as they depend on the unknown statistical properties of the variable process under study. Therefore, the sampled autocovariance functions cannot be easily fitted by models.
The most frequently used estimator of the PSD is called the “periodogram”, and its statistical properties are much superior. If computed over a certain range of frequencies, the periodogram estimates are asymptotically independent random variables with a known distribution. Their error depends on the intrinsic PSD (which is unknown), but this deficit can be overcome by considering either the logarithm of the periodogram [in which case the log-periodogram error is know, i.e. 229], or by “smoothing” the periodogram estimates. The periodogram is an asymptotically unbiased estimate of the intrinsic PSD, although in the case of “red-noise” power spectra the lowest frequency estimates can be biased (this is the so-called “red-leakage” bias, see 229, and references therein). Given the superior statistical properties of the PSD estimates and the fact that it can help us easily detect characteristic timescales (either signaling periodicity or otherwise), PSD analysis has been the dominant method of studying the variability in the observed X-ray light curves of AGN.
@@ 3.3 Results from the PSD analysis of the X-ray light curves
Another interesting result from the power spectrum analysis of the EXOSAT light curves was that the ME PSDs were significantly flatter than the lower-energy PSDs [230]. This is in contradiction with the expectations from inverse Compton scattering of low-energy photons in a single-temperature X-ray corona. Hard X-ray photons should undergo more scatterings in this case, hence high frequency variations should be washed out, and the PSD at high energies should be separately steeper, contrary to what is observed. Similar results were also reported for galactic X-ray black hole binaries (GBHBs; e.g. Nowak et al. 215).
@@ 3.3.1 The RXTE/XMM-Newton legacy
It was long proposed that the X-ray PSDs of AGN cannot continue with a steep slope to very low frequencies because the total variance would then be unphysically large. There were indications that PSD “breaks” to a flatter slope at frequencies below a characteristic “break” or “bending” frequency () in the PSD analysis of past observations of AGN [e.g. 231]. However, the first clear and unambiguous detection of such a “break timescale” was based on the PSD analysis of the RXTE light curve of NGC 3516 and was reported by [80]. Soon after, [303] and [186] published the results of a detailed study of the power spectrum of a few Seyferts using RXTE light curves and reported further detections of such bending frequencies. Based on the results of these papers, a picture of a power spectrum with a slope of at low frequencies that steepens to above emerged in the 2–10 keV band.
The next significant improvement in the determination of the AGN X-ray PSDs was made possible by the combination of data from RXTE and XMM-Newton in order to calculate the power spectrum over a very wide frequency range. Arguably, the best example of such a well-defined PSD is the power spectrum of NGC 4051 presented by McHardy et al. [201, see the left panel in Fig. 26]. This power spectrum is well fitted by a “bending power law (BPL) model” of the form:
where and are the PSD slopes below and above , respectively, while is equal to PSD(. In the case where , then PSD. At frequencies much lower than , then PSD and, therefore, is representative of the PSD amplitude (when ).
@@
There is plenty of observational evidence on the increase in the characteristic frequencies in the GBHB X-ray power spectra with increasing X-ray flux, that is, with an increasing accretion rate in this case (e.g. Axelsson et al. 4 for Cyg X-1, Reig et al. 257 for GRO J1655-40, Alabarta et al. 3 for MAXI J1727–203). Assuming that the X-ray variability properties of AGN and GBHBs are similar, this observational evidence supports the findings of [195]. If true, this result puts interesting constraints on the physical timescale that could be responsible for in AGN.
@@
It is possible that the corona is highly magnetised and is powered by magnetic processes. For example, according to Haardt et al. [113, see also ] the energy that powers the X-ray corona may be stored in magnetic field structures that produce active X-ray regions on top of the disc. The energy is stored in the magnetic field on a charge timescale, , and is released on a shorter discharge timescale, . In this scenario, the main characteristic timescales of the X-ray corona are and . The charge timescale is given by:
The left panel of Fig. 27 shows a plot of log( versus BH mass, for a small sample of AGN. We used data from [104] for NGC 4395 and [198] for Mrk 766, who studied the power spectra using only XMM-Newton light curves. For all the other objects, we used results from power spectral analysis of combined RXTE and XMM-Newton PSDs (McHardy et al. 201 for NGC 4051, McHardy et al. 189 for MCG -6-30-15, Uttley and McHardy 302 for NGC 3227, Markowitz 175 for NGC 7469, Summons et al. 260 for NGC 3783, and Markowitz et al. 186 for NGC 4151, NGC 3516, NGC 5548 and Fairall 9). For the BH mass estimates, we used the measurements listed in the “AGN Black Hole Mass Database”, which have been calculated by spectroscopic reverberation mapping studies [assuming a mean factor of 4.3, 30]. Solid, dashed, and dotted lines show the theoretical MBH relation assuming that , in the case where and 0.01, respectively. Clearly, the hypothesis that can explain most of the scatter in the observed relation between and MBH, as long as the accretion rate decreases from the smaller to the larger BH mass objects, which is what is observed in low redshift samples of AGN, such as the BASS sample [e.g. 153].
@@
The right panel in Fig. 27 shows the product PSD( for Seyeferts where this product can be calculated using the best-fit results from the BPL model fits to the 2-10 keV PSDs. This quantity is representative of the PSD amplitude (as we explained in §3.3.1). We have used the results of [201, 189, 302] for NGC 4051, MCG-6-30-15, and NGC 3227, respectively, and [186] for Fairall 9, NGC 5548 and NGC 3516. These authors used RXTE and XMM-Newton light curves to compute the power spectrum over a wide range of frequencies. For the rest of the sources (i.e. Mrk 335, Mrk 766, NGC 4395 and NGC 6860) we used the [104] results. PSD amplitudes are plotted as a function of log( (we used the Gupta et al. 101 measurements of to calculate ).
@@
The number of direct measurements of the PSD amplitude so far has been rather small, and we cannot draw any firm conclusions. The data so far indicate a distribution of PSD amplitudes with the same mean for all AGN, but further PSD analysis of combined RXTE and XMM light curves is necessary to reach conclusive results.
@@
The PSD slope at frequencies below is generally believed to be very close to . This is a rather well established fact, based on the results of the bending PL model fits to PSDs computed using long-term RXTE light curves. The left panel of Fig. 28 shows the best fit for NGC 4051 [201], MCG-6-30-15 [189], Fairall 9, NGC5548, NGC3516 and NGC4151 [186], as well as NGC3227 [302], NGC3783 [260] and NGC7469 [175]. The solid horizontal line indicates the value of . Clearly, all the low-frequency PSD slope estimates are consistent with this value (within the error). The PSDs of NGC4051 and MCG-6-30-15 (201 and 189, respectively) are particularly important in this respect, as they show that the slope of -1 extends over and orders of magnitude below . This is typical of the XRB PSDs when they are in their so-called “soft state”.